Construct a 95% confidence interval for the true proportion of all employees of Handy Glass Pty who are members of the company’s medical aid scheme, given that 75 out of a sample of 135 employees are members of the company’s medical aid scheme. 


1
Expert's answer
2020-10-20T18:06:03-0400

"\text{The confidence interval for the }\\
\text{proportion can be calculated as follows:}\\
 \hat p ±Z_{\frac{α}{2}}(\sqrt {\frac{\hat p (1-\hat p )}{n}}),\\
 \hat p =\frac{75}{135}\approx0.56,\\
 Z_{\frac{α}{2}}=Z_{\frac{0.05}{2}}=Z_{0.025}=1.96,\\
 \text{the lower limit }= 0.56-1.96(\sqrt {\frac{0.56 (1-0.56 )}{135}})\\
≈0.476,\\
 \text{the upper limit }= 0.56+1.96(\sqrt {\frac{0.56(1-0.56)}{135}})\\
≈0.644,\\
 \text{so the confidence interval is}\\
0. 476\leq \ p \leq 0.644"